OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line containing the diagonal of a square whose vertices are A(-3, 3), B(3, 3), C(3, -3), and D(-3, -3). Find two equations, one for each diagonal.

OpenStudy (anonymous):

ok start by finding slope using the formula (y2-y1) / (x2-x1)

OpenStudy (anonymous):

do i find the slope between each of the vertices?

OpenStudy (anonymous):

exactly....

OpenStudy (anonymous):

i got 0/6 i feel like that isn't right.

OpenStudy (anonymous):

substituting A (-3,3) for (x1,y1) and C (3,-3) for (x2,y2) you get, (3-(-3)) / (-3-3) = 6/(-6) = -1 so the slope of the first line is -1

OpenStudy (anonymous):

now use this formula, (y-y1) = m (x-x1) where (x1,y1) is a point on the line and "m" is the slope. (y-3) = (-1) * (x-(-3)) y-3 = -1 (x+3) y-3 = -x-3

OpenStudy (anonymous):

you will get x+y = 0 . similarly do for rest second line (B to D). first find the slope again using the same formula. you do this by using the formula (y2-y1) / (x2-x1) where (x1,y1) is one point on the line and (x2,y2) is another point on the line. so substituting B (3,3) for (x1,y1) and D (-3,-3) for (x2,y2) you get: (-3-3) / (-3-3) = (-6) / (-6) = 1 so the slope of the second line is 1

OpenStudy (anonymous):

(y-y1) = m (x-x1) so using point B (3,3) for (x1,y1) and -1 for "m" you get: (y-3) = 1* (x-3) y-3 = x-3 -x + y = 0 It's DONE ;) @aleishealarson

OpenStudy (anonymous):

x+y = 0 and -x + y = 0 Hope it helps a bit ^_^

OpenStudy (anonymous):

thank you so much!!! xxx

OpenStudy (anonymous):

always a pleasure :) :P @aleishealarson